OpenStudy (anonymous):

Find the surface area of the ball. Leave the answer in terms of pi .d=68mm @Directrix

Directrix (directrix):

A ball is a sphere in this problem.

OpenStudy (anonymous):

yea the picture was a tennis ball

Directrix (directrix):

Surface Area of Sphere If the diameter is 68, then the radius is 34 A = 4 * π * r² A = 4 * π * 34²

OpenStudy (anonymous):

do i use the pi in this one

Directrix (directrix):

Keep pi as pi. Do not approximate.

OpenStudy (anonymous):

how exactly do i do that

Directrix (directrix):

I got 4624π square mm

OpenStudy (anonymous):

i got it

Directrix (directrix):

Good, next question.